# Exploring the role of accretion disk geometry in shaping black hole shadows

### v1

@@ Abstract
We study black hole imaging in the context of geometrically thick accretion disks in Schwarzschild spacetime. By decomposing the emitting region into a set of one-dimensional luminous segments, each characterized by its inclination angle and inner radius, we construct transfer functions that capture key image features-namely, the direct image, lensing ring, and photon ring. This approach allows a unified treatment of disk geometry and viewing angle. We explore three regimes: optically thin, optically thick, and partially optically thick disks. For optically thin flows, increasing the disk thickness broadens the lensing ring, gradually bridging the photon ring and the direct image. The photon ring remains narrow, but its position robustly defines the innermost edge of the lensing structure. In the optically thick case, image features are primarily determined by the first intersection of traced light rays with the disk, and we provide analytical criteria for the presence of lensing and photon rings based on the critical deflection angles. For partially optically thick disks, we adopt a simplified radiative transport model and find a critical absorption coefficient $\chi\sim(6M\psi_{0})^{-1}$ beyond which the image rapidly transitions from an optically thin- to thick-disk appearance. These results help clarify the respective roles of the photon and lensing rings across different disk configurations, and may offer a useful framework for interpreting future high-resolution black hole observations.
@@ 1 Introduction
Since the Event Horizon Telescope (EHT) successfully imaged the shadows of the supermassive black holes in M87* EventHorizonTelescope:2019dse ; EventHorizonTelescope:2019ggy and Sagittarius A* (Sgr A*) EventHorizonTelescope:2022wkp , black hole shadow observations have become a cornerstone of modern astrophysics, offering a unique opportunity to test general relativity in the strong-field regime and to probe the nature of black holes. Interpreting such images critically depends not only on the background spacetime geometry but also on the physical processes occurring near the black hole-particularly the structure and geometry of the surrounding accretion flow.
It is well understood that the spacetime geometry, which governs the trajectories of photons via gravitational curvature, plays a crucial role in shaping the black hole shadow. However, for a fixed spacetime background, the optical appearance of the black hole can be substantially affected by the structure and dynamics of the surrounding accretion flow. Different accretion geometries-such as geometrically thin disks, thick disks, and spherically symmetric flows-can produce distinct observational signatures. This naturally raises the question: between the spacetime geometry and the accretion flow, which plays the dominant role in shaping the black hole shadow?
Recent studies focusing on Schwarzschild black holes have presented differing perspectives on this issue. Gralla et al. Gralla:2019xty argue that the structure of the accretion flow plays a more dominant role in shaping the observed black hole shadow than intrinsic features such as the photon ring. In contrast, Narayan et al. Narayan:2019imo demonstrate that the shadow edge is largely governed by the underlying spacetime geometry, with minimal dependence on the specific structure of the accreting material. These differing conclusions stem from the distinct accretion models considered: Gralla et al. mainly investigate a thin disk configuration, whereas Narayan et al. focus on spherically symmetric, optically thin flows-specifically, both static gas and infalling gas, with the latter modeled as a Bondi-type inflow Bondi:1952 ; Abramowicz:2011xu .
@@
Figure 1: Schematic meridional cross-sectional view ($x-z$ plane) of a geometrically thick accretion disk. The disk is symmetric about the equatorial plane and extends outward from an inner radius $r_{\mathrm{in}}$. In general, the relative thickness is characterized by $h/r$. In this paper, we consider a simplified scenario where the thickness is described by an half-opening angle $\psi_{0}$ at $r_{\mathrm{in}}$.
@@ 2.1 Coordinate systems
$d_{P^{\prime}}=r_{\rm obs}\tan\beta\,,$ (3)
where $\beta$ is defined as the angle between the observer-black hole axis and the incident photon trajectory. For a distant observer with $r_{\rm obs}\gg M$, it can be shown that
@@ 2.2.1 Disk geometry and dynamics of the flow
  • Inner edge $r_{\mathrm{in}}$ and half-opening angle $\psi_{0}$  [^1] The half-opening angle $\psi_{0}$ (or the opening angle $2\psi_{0}$) is a commonly used parameter in the literature Abramowicz:2011xu ; Arshakian:2004kh to characterize the thickness of accretion flows.. The accretion disk is symmetric about the equatorial plane and axisymmetric around the $z$-axis, extending outward from an inner radius $r_{\mathrm{in}}$. The vertical thickness of the disk is characterized by an half-opening angle $\psi_{0}$ at $r_{\mathrm{in}}$, see Fig. 1. We consider the following parameter ranges: $r_{\rm h}<r_{\mathrm{in}}\leq r_{\mathrm{ISCO}}$ and $20<\psi_{0}\leq\pi/$ 2.
@@
      We consider deviations from pure free fall which are modeled by $v^{r}=\kappa_{\mathrm{ff}}v^{r}_{\mathrm{ff}}$ with $10\leq\kappa_{\mathrm{ff}}<$ 1. A representative example is the ADAF, where the flow is described by a specific heat ratio of 5/3 Narayan:1994xi ; kato2008black ; Abramowicz:2011xu .
@@ 2.2.2 Emission profile and absorption
For an optically thick disk, the local emitted flux-that is, the radiative energy emitted per unit time per unit area of the disk surface-is modeled as
@@
$\delta F_{\rm em}\approx\frac{\epsilon\dot{M}}{4\pi\psi_{0}}\left(-\Omega+\frac{r_{\rm in}^{2}\,\Omega(r_{\rm in})}{r^{2}}\right)\frac{\mathrm{d}\Omega}{\mathrm{d}r}\sqrt{B}\,\delta l,$ (9)
where $\psi_{0}$ is the half-opening angle defined in Sec. 2.2.1. The factor $\sqrt{B}$ in (9) accounts for the proper radial distance when computing the volume element. The interval $\delta l$ could be expressed as
$\delta l=\left(B-B^{2}b^{2}/r^{2}\right)^{-1/2}{\delta r}$ (10)
via the $r-$component of the photon’s 4-momentum. At locations where $\delta r$ vanishes, an alternative expression can be obtained using the $\phi-$component of the photon’s 4-momentum.
@@
$\frac{k_{r}}{k_{t}}$ =±$=\pm\frac{\sqrt{1-B\,b^{2}/r^{2}}}{B},$ (13e)
@@ 3 Imaging of One-dimensional Segments
Figure 3: Sketch of one-dimensional luminous segments lying in the photon orbital plane. The colorful thick solid lines represent the emitting segments. Thin black lines illustrate representative photon trajectories, while the dashed line indicates the observer-black hole axis.
A schematic illustration of luminous segments with different inclination angles is presented in Fig. 3. Here, $\varphi_{0}\in(0,\pi]$ denotes the inclination angle of an individual luminous segment, which should not be confused with $\theta_{0}$, the inclination angle of the entire accretion disk as introduced in Sec. 2. The latter follows the standard convention commonly used in astrophysics. Without loss of generality, we restrict our analysis in this section to the image line corresponding to photons with positive impact parameters. Under this setup, a thin accretion disk with an inclination angle $\theta_{0}$ can be mathematically modeled as a continuous distribution of radial luminous segments, each characterized by an inclination angle $\varphi_{0}$ ranging from $[\pi/2-\theta_{0},,\pi/2+\theta_{0}]$.
@@ 3.1 Transfer functions for one-dimensional segments
Fig. 4 illustrates the first three transfer functions for luminous segments at different inclination angles. An important features can be identified from the figure is that, for the second and third transfer functions, a critical transition occurs near $r\approx r_{\rm{sp}}=3M$, where the impact parameter $b$ attains $3\sqrt{3}M$ and is nearly independent of the inclination angle $\varphi_{0}$. This radius corresponds to the photon sphere and plays a key role in shaping the image structure. For radii larger than $r_{\rm{sp}}$, the transfer functions exhibit approximately monotonic behavior: the impact parameter $b$ tends to decrease as the inclination angle $\varphi_{0}$ increases. In contrast, for $r<r_{\rm{sp}}$, the trend reverses: $b$ increases with increasing $\varphi_{0}$. The first transfer function does not exhibit this monotonic behavior for radii greater than $r_{\rm{sp}}$. In the notation of Ref. Gralla:2019xty , the image components associated with the first three transfer functions are referred to as the “direct image,” “lensing ring,” and “photon ring,” respectively. It should also be noted that for segments with inner edge $r_{\rm in}>r_{\rm sp}$, the appearance of a lensing ring or photon ring does not necessarily require multiple intersections with the segments. This distinction becomes particularly important in the optically thick case, where light rays typically intersect the emitting surface only once. Therefore, the classical definitions-based on multiple intersections with the disk-no longer strictly apply. In this work, we adopt the following generalized definitions:
  • Lensing ring: the set of image points corresponding to light rays whose total deflection angle exceeds $\pi$.
  • Photon ring: the set of image points corresponding to light rays whose total deflection angle exceeds $2\pi$.
@@ 3.2 Image of one-dimensional segments
One important application of the transfer function is to determine the inner boundary of the image line, as well as the minimum radius at which potential peaks-such as those corresponding to the lensing ring or photon ring-can appear. An explicit example of this can be seen by comparing the transfer functions in the top-right panel of Fig. 4 with the flux profiles shown in Fig. 5.444In calculating the flux contribution from each luminous segment, we continue to model it as a one-dimensional object, but assign it an infinitesimal area element to ensure a physically meaningful emission flux.
Specifically, in the top-right panel of Fig. 4, the intersection points between the transfer functions and the horizontal line $r=6M$ indicate the inner edge of the image line or the minimum radius at which peaks may appear. These features are directly reflected in the flux distribution shown in Fig. 5. In the optically thin regime, the correspondence is typically one-to-one, with each intersection point mapping directly to the inner edge either of the image line or a distinct brightness peak. In contrast, for optically thick emission, the presence of peaks associated with the lensing ring and photon ring depends sensitively on the inclination angle $\varphi_{0}$ and the segment’s inner radius $r_{\rm in}$; such peaks may be absent in some configurations-for instance, see the case with $\varphi_{0}=40^{\circ}$ in the right panel of Fig. 5.
In fact, an analytical condition can be derived for the appearance of the photon ring in the optically thick case. The critical photon orbit-i.e., a null geodesic arriving from infinity with the critical impact parameter $b=3\sqrt{3}M$-asymptotically spirals toward the photon sphere. The radial trajectory of this orbit as a function of the deflection angle $\varphi$ is given by the following expression Chandrasekhar1985 :
@@
Another observation from the left panel of Fig. 5 is that, for $\varphi_{0}=90^{\circ}$ and $140^{\circ}$, the observed flux profile contains three distinct peaks-corresponding to the photon ring, lensing ring, and direct image-though the visibility of the photon ring peak may depend on the experimental resolution. In contrast, for $\varphi_{0}=40^{\circ}$, only two peaks are present: the photon ring and the lensing ring. This is because the inner edge of the direct image extends into the region inside the critical impact parameter, and its flux contribution is spread over a wider area, resulting in a smoother profile that lacks a clearly defined peak due to the overlapping influence of the photon and lensing rings. In the right panel of Fig. 5, which shows the optically thick scenario, we observe a broad and smooth flux enhancement for $\varphi_{0}=40^{\circ}$ associated with the direct image. Since the photon ring and lensing ring are absent in this case, the direct image becomes more apparent, albeit without a sharply defined peak.
@@ 4 Imaging of Accretion Disks
Moreover, we classify our results by opacity-into optically thin, optically thick, and partially optically thick regimes-to highlight the distinct radiative features of each. In summary, within the scope allowed by our computational resources, we have incorporated a wide range of parameters in an effort to gain a more comprehensive understanding of black hole imaging.

### v4

@@ Abstract
We study black hole imaging in the context of geometrically thick accretion disks in Schwarzschild spacetime. By decomposing the emitting region into a set of one-dimensional luminous segments, each characterized by its inclination angle and inner radius, we construct transfer functions that capture key image features—namely, the direct image, lensing ring, and photon ring. This approach allows a unified treatment of disk geometry and viewing angle. We explore three regimes: optically thin, optically thick, and partially optically thick disks. For optically thin flows, increasing the disk thickness (characterized by the half-opening angle $\psi_{0}$) broadens the lensing ring, gradually bridging the photon ring and the direct image. The photon ring remains narrow, but its position robustly defines the innermost edge of the lensing structure. In the optically thick case, image features are primarily determined by the first intersection of traced light rays with the disk, and we provide analytical criteria for the presence of lensing and photon rings based on the critical deflection angles. For partially optically thick disks, we adopt a simplified radiative transport model and find a critical absorption coefficient $\chi\sim(6M\psi_{0})^{-1}$ beyond which the image rapidly transitions from an optically thin- to thick-disk appearance. These results help clarify the respective roles of the photon and lensing rings across different disk configurations, and may offer a useful framework for interpreting future high-resolution black hole observations.
arxiv: 2506.21148
@@ 1 Introduction
Since the Event Horizon Telescope (EHT) successfully imaged the shadows of the supermassive black holes in M87* EventHorizonTelescope:2019dse ; EventHorizonTelescope:2019ggy and Sagittarius A* (Sgr A*) EventHorizonTelescope:2022wkp , black hole shadow observations have become a cornerstone of modern astrophysics, offering a unique opportunity to test general relativity in the strong-field regime and to probe the nature of black holes Vagnozzi:2022moj ; Virbhadra:1999nm ; Adler:2022qtb ; Zhang:2024lsf ; Sokoliuk:2022owk ; Virbhadra:2022iiy ; Hou:2023bep ; Li:2025ixk ; Sui:2023yay ; Li:2024oyc . Interpreting such images critically depends not only on the background spacetime geometry but also on the physical processes occurring near the black hole—particularly the structure and geometry of the surrounding accretion flow.
It is well understood that the spacetime geometry, which governs the trajectories of photons via gravitational curvature, plays a crucial role in shaping the black hole shadow. However, for a fixed spacetime background, the optical appearance of the black hole can be substantially affected by the structure and dynamics of the surrounding accretion flow. Different accretion geometries—such as geometrically thin disks, thick disks, and spherically symmetric flows—can produce distinct observational signatures. This naturally raises the question: between the spacetime geometry and the accretion flow, which plays the dominant role in shaping the black hole shadow?
Recent studies focusing on Schwarzschild black holes have presented differing perspectives on this issue. Gralla et al. Gralla:2019xty argue that the structure of the accretion flow plays a more dominant role in shaping the observed black hole shadow than intrinsic features such as the photon ring. In contrast, Narayan et al. Narayan:2019imo demonstrate that the shadow edge is largely governed by the underlying spacetime geometry, with minimal dependence on the specific structure of the accreting material. These differing conclusions stem from the distinct accretion models considered: Gralla et al. mainly investigate a thin disk configuration, whereas Narayan et al. focus on spherically symmetric, optically thin flows—specifically, both static gas and infalling gas, with the latter modeled as a Bondi—type inflow Bondi:1952 ; Abramowicz:2011xu .
@@
Figure 1: Schematic meridional cross-sectional view (xzx-z plane) of a geometrically thick accretion disk. The disk is symmetric about the equatorial plane and extends outward from an inner radius $r_{\mathrm{in}}$. In general, the relative thickness is characterized by $h/r$. In this paper, we consider a simplified scenario where the thickness is described by an half-opening angle $\psi_{0}$ at $r_{\mathrm{in}}$.
nai
@@ 2.1 Coordinate systems
$d_{P^{\prime}}\simeq r_{\rm obs}\tan\beta\,$, (3)
where $\beta$ is defined as the angle between the observer-black hole axis and the incident photon trajectory. Eq. (3) holds approximately under the assumption that the observer is located far from the black hole, i.e., $r_{\rm obs}\gg M$, so that the local geometry can be treated as nearly flat and the angle $\beta$ remains sufficiently small. In this regime, $\beta$ can be identified with the angular radius commonly used in astronomy. For a distant observer, it can be shown that
@@ 2.2.1 Disk geometry and dynamics of the flow
  • Inner edge $r_{\mathrm{in}}$ and half-opening angle $\psi_{0}$  [^1] The half-opening angle $\psi_{0}$ (or the opening angle $2\psi_{0}$) is a commonly used parameter in the literature Abramowicz:2011xu ; Arshakian:2004kh to characterize the thickness of accretion flows.. The accretion disk is symmetric about the equatorial plane and axisymmetric around the $z$-axis, extending outward from an inner radius $r_{\mathrm{in}}$. The vertical thickness of the disk is characterized by an half-opening angle $\psi_{0}$ at $r_{\mathrm{in}}$, see Fig. 1. We consider the following parameter ranges: $r_{\rm h}<r_{\mathrm{in}}\leq r_{\mathrm{ISCO}}$ and $0<\psi_{0}\leq\pi/2$.
@@
      We consider deviations from pure free fall which are modeled by $v^{r}=\kappa_{\mathrm{ff}}v^{r}_{\mathrm{ff}}$ with $0\leq\kappa_{\mathrm{ff}}<1$. A representative example is the ADAF, where the flow is described by a specific heat ratio of 5/3 Narayan:1994xi ; kato2008black ; Abramowicz:2011xu .
@@ 2.2.2 Emission profile and absorption
For an optically thick disk, the local emitted flux—that is, the radiative energy emitted per unit time per unit area of the disk surface—is modeled as
@@
$\delta F_{\rm em}=q^{-}\,\sqrt{B}\,\delta l\approx\frac{\epsilon\dot{M}}{4\pi\psi_{0}}\left(-\Omega+\frac{r_{\rm in}^{2}\,\Omega(r_{\rm in})}{r^{2}}\right)\frac{\mathrm{d}\Omega}{\mathrm{d}r}\sqrt{B}\,\delta l,$ (9)
where $q^{-}$ denotes the rate of radiative energy loss per coordinate volume element (a derivation of $q^{-}$ is provided in Appendix B) [^2] The derivation of $q^{-}$ in Appendix B is purely classical, where the volume element is implicitly assumed to be $r^{2}\sin\theta{\rm d}r{\rm d}\theta{\rm d}\phi$. To convert this coordinate volume element into the proper volume element, a factor $\sqrt{B}$ should be included. For a more relativistic treatment, one may adopt the Novikov-Thorne model novikov1973astrophysics , which, however, involves complicated calculations, especially when generalized to geometrically thick disks. Given the goals of this paper, our simplified approach suffices to capture the essential physics., and $\psi_{0}$ is the half-opening angle defined in Sec. 2.2.1 The factor $\sqrt{B}$ in (9) accounts for the proper radial distance, converting the coordinate volume element into the proper volume element. The interval $\delta l$ represents the spatial distance travelled by a photon with four-momentum $k^{\mu}$, as measured by a local observer with four-velocity $V^{\mu}$. It can be expressed as
$\delta l=\sqrt{(g_{\mu\nu}+V_{\mu}V_{\nu})k^{\mu}k^{\nu}}\,\delta\lambda=-k_{\mu}V^{\mu}\,\delta\lambda\,.$ (10)
In numerical computations, the affine parameter interval $\delta\lambda$ can be related to the $r$-component of the photon’s four-momentum, yielding
$\delta l=\left(B-B^{2}b^{2}/r^{2}\right)^{-1/2}{\delta r}\,$, (11)
where the four-velocity $V^{\mu}$ of a local static observer has been used. (Note that this local static observer is hypothetical and is assumed to reside on (or within) the accretion disk for illustrative purposes. It should not be confused with the distant observer.) At locations where $\delta r$ vanishes (e.g., near turning points), an alternative expression involving the $\phi$-component of the photon’s four-momentum may be used.
@@
$\frac{k_{r}}{k_{t}}$ $=\pm\frac{\sqrt{1-B\,b^{2}/r^{2}}}{B},$ (14e)
@@ 3 Imaging of One-dimensional Segments
Figure 3: Sketch of one-dimensional luminous segments lying in the photon’s orbital plane. The colorful thick solid lines depict representative emitting segments. Thin black lines illustrate representative photon trajectories, while the dashed line indicates the observer-black hole axis.
A schematic illustration of luminous segments with different inclination angles is presented in Fig. 3. Here, $\varphi_{0}\in(0,\pi]$ denotes the inclination angle of an individual luminous segment, which should not be confused with $\theta_{0}$, the inclination angle of the entire accretion disk as introduced in Sec. 2. The latter follows the standard convention commonly used in astrophysics. As illustrated in Fig. 3, regardless of the inclination angle of the accretion disk, we assume that the luminous segments lie in the photon’s orbital plane rather than the physical disk plane. It should be emphasized that a two- or three-dimensional accretion disk can always be viewed as a composition of such luminous segments, regardless of its inclination angle. Without loss of generality, we restrict our analysis in this section to the image line corresponding to photons with positive impact parameters. Under this setup, a thin accretion disk with an inclination angle $\theta_{0}$ can be mathematically modeled as a continuous distribution of radial luminous segments, each characterized by an inclination angle $\varphi_{0}$ ranging from $\pi/2-\theta_{0},\,\pi/2+\theta_{0}]$.
@@ 3.1 Transfer functions for one-dimensional segments
Fig. 4 illustrates the first three transfer functions for luminous segments at different inclination angles. An important features can be identified from the figure is that, for the second and third transfer functions, a critical transition occurs near $r\approx r_{\rm{sp}}=3M$, where the impact parameter $b$ attains $3\sqrt{3}M$ and is nearly independent of the inclination angle $\varphi_{0}$. This radius corresponds to the photon sphere and plays a key role in shaping the image structure. For radii larger than $r_{\rm{sp}}$, the transfer functions exhibit approximately monotonic behavior: the impact parameter $b$ tends to decrease as the inclination angle $\varphi_{0}$ increases. In contrast, for $r<r_{\rm{sp}}$, the trend reverses: $b$ increases with increasing $\varphi_{0}$. The first transfer function does not exhibit this monotonic behavior for radii greater than $r_{\rm{sp}}$. In the notation of Ref. Gralla:2019xty , the image components associated with the first three transfer functions are referred to as the “direct image,” “lensing ring,” and “photon ring,” respectively. It should also be noted that for segments with inner edge $r_{\rm in}>r_{\rm sp}$, the appearance of a lensing ring or photon ring does not necessarily require multiple intersections with the segments. This distinction becomes particularly important in the optically thick case, where light rays typically intersect the emitting surface only once. Therefore, the classical definitions—based on multiple intersections with the disk—no longer strictly apply. In this work, we adopt the following generalized definitions:
  • Lensing ring: the set of image points formed by light rays whose total azimuthal angle change (in the orbital plane) outside the horizon exceeds $\pi$.
  • Photon ring: the set of image points formed by light rays whose total azimuthal angle change (in the orbital plane) outside the horizon exceeds $2\pi$.
@@ 3.2 Image of one-dimensional segments
One important application of the transfer function is to determine the inner boundary of the image line, as well as the minimum radius at which potential peaks—such as those corresponding to the lensing ring or photon ring—can appear. An explicit example of this can be seen by comparing the transfer functions in the top-right panel of Fig. 4 with the flux profiles shown in Fig. 5.555In calculating the flux contribution from each luminous segment, we continue to model it as a one-dimensional object, but assign it an infinitesimal area element to ensure a physically meaningful emission flux.
Specifically, in the top-right panel of Fig. 4, the intersection points between the transfer functions and the horizontal line $r=6M$ indicate the inner edge of the image line or the minimum radius at which peaks may appear. These features are directly reflected in the flux distribution shown in Fig. 5. In the optically thin regime, the correspondence is typically one-to-one, with each intersection point mapping directly to the inner edge either of the image line or a distinct brightness peak. In contrast, for optically thick emission, the presence of peaks associated with the lensing ring and photon ring depends sensitively on the inclination angle $\varphi_{0}$ and the segment’s inner radius $r_{\rm in}$; such peaks may be absent in some configurations—for instance, see the case with $\varphi_{0}=40^{\circ}$ in the right panel of Fig. 5.
In fact, an analytical condition can be derived for the appearance of the photon ring in the optically thick case. The critical photon orbit—i.e., a null geodesic arriving from infinity with the critical impact parameter $b=3\sqrt{3}M$—asymptotically spirals toward the photon sphere. The radial trajectory of this orbit as a function of the deflection angle $\varphi$ is given by the following expression Chandrasekhar1985 :
@@
Another observation from the left panel of Fig. 5 is that, for $\varphi_{0}=90^{\circ}$ and $140^{\circ}$, the observed flux profile contains three distinct peaks—corresponding to the photon ring, lensing ring, and direct image—though the visibility of the photon ring peak may depend on the experimental resolution. In contrast, for $\varphi_{0}=40^{\circ}$, only two peaks are present: the photon ring and the lensing ring. This is because the inner edge of the direct image extends into the region inside the critical impact parameter, and its flux contribution is spread over a wider area, resulting in a smoother profile that lacks a clearly defined peak due to the overlapping influence of the photon and lensing rings. In the right panel of Fig. 5, which shows the optically thick scenario, we observe a broad and smooth flux enhancement for $\varphi_{0}=40^{\circ}$ associated with the direct image. Since the photon ring and lensing ring are absent in this case, the direct image becomes more apparent, albeit without a sharply defined peak.
@@ 4 Imaging of Accretion Disks
Moreover, we classify our results by opacity—into optically thin, optically thick, and partially optically thick regimes—to highlight the distinct radiative features of each. In summary, within the scope allowed by our computational resources, we have incorporated a wide range of parameters in an effort to gain a more comprehensive understanding of black hole imaging.